OpenStudy (tinyclown):

find the antiderivative of x^2/((2x^3-12)^4) using substitution

OpenStudy (tinyclown):

now what i got was -1/(24(2x^3-12)^3)) + C

OpenStudy (tinyclown):

and to check, i did the antiderivative on wolfgram and got -1/(144(x^3-6)^3)) or something i don t remember

OpenStudy (tinyclown):

i graphed both of those and they were the same function but

OpenStudy (tinyclown):

neither is an answer to the quiz question (it s multiple choice)...now there is an option that says none of the above but also says that the problem can t be solved through substitution despite the fact that s exactly how i solved it

zepdrix (zepdrix):

Hmm maybe your coefficient is a little off.. I'm coming up with -1/18 as opposed to your -1/24

OpenStudy (tinyclown):

okay so that s actually one of the options let me recheck my math

zepdrix (zepdrix):

\[\large\rm \int\limits\frac{x^2dx}{(\color{orangered}{2x^3-12})^4}\] \[\large\rm \color{orangered}{u=2x^3-12},\qquad\qquad du=6x^2dx,\qquad\qquad\frac{1}{6}du=x^2dx\]

zepdrix (zepdrix):

Yah maybe something got mixed up with your power rule :) You should get a -1/3 when integrating, not a -1/4.

OpenStudy (tinyclown):

oh yeah yeah that s right !! thank you so much !!

zepdrix (zepdrix):

np